OpenStudy (anonymous):

An isosceles trapezoid has measurements as follows. What is the area of the entire trapezoid? upper base is 10, height is 12, slant height is 13, and the longer part of the lower base is 14

OpenStudy (anonymous):

what do you mean by the 'longer part of the lower base'?

OpenStudy (anonymous):

ill show u a pic

OpenStudy (anonymous):

ill show u a pic

OpenStudy (anonymous):

ok shortest part of lower base = sqrt(13^2 - 12^2) = 5 length of lower base = 19 area = 12/2 * (19+10) = 6* 29 = 174